# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 6.12 External Validity for Minimalist Market Design: School Choice Reforms in England and Chicago
Following the reform at BPS, and two subsequent reforms in Chicago and England that exhibited strong parallels, we made the following point in Pathak and Sönmez (2013):
@@
A defining characteristic of minimalist market design is its focus on stakeholder objectives. It begins by identifying the desiderata that matter most to stakeholders, then uncovers the failures of existing institutions in light of these goals. This objective-driven approach stems from my thesis that stakeholders may only consider alternative, better institutions once this advocacy stage has been reached. Let’s refer to this stage in a potential redesign as the initial phase of the reform path.
Following Abdulkadiroğlu and Sönmez (2003), in Boston, officials explored alternatives to the Boston mechanism only after discovering its vulnerability to preference manipulation, marking the initial phase of the reform path. Subsequently, they selected a strategy-proof mechanism, as also suggested in Abdulkadiroğlu and Sönmez (2003), representing the second stage of the reform path. From my perspective, this two-stage reform path under research following minimalist market design not only helps uncover mistakes in original designs but also mimics the intended natural evolution of mechanisms through its remedial proposals.
Building on our analysis in Pathak and Sönmez (2013), I will now examine two subsequent reforms in Chicago and England that followed the same reform path, albeit without any involvement from design economists. This notable parallel lends external validity not only to the policy guidance of Abdulkadiroğlu and Sönmez (2003) but also to minimalist market design itself, supporting my view that this design framework “mimics” the “intended” natural evolution of target institutions.
@@
In England, the nationwide 2003 School Admissions Code mandated that local authorities, an operating body similar to a U.S. school district, should coordinate student admissions through a centralized mechanism. With the 2003 code, two classes of mechanisms were recommended.
The first of these two classes was the DA and its “capped” versions, where the maximum number of schools that can be ranked in student preferences is limited to a fixed number. The second one was a generalization of the Boston mechanism called a first preference first system and its capped versions. However, four years later, all versions of the first preference first systems, including the Boston mechanism, were banned throughout England with the 2007 School Admissions Code. As a result, all local authorities in England were using variants of DA by 2007.
@@
Thus, much like in the 2005 school choice reform at BPS, the vulnerability of a variant of the Boston mechanism to strategic manipulation and related concerns led to its demise in England as well. For our purposes, the most significant aspect is the reform path underlying the 2007 school choice reform in England, which closely parallels the 2005 BPS reform and aligns with the policy guidance in Abdulkadiroğlu and Sönmez (2003). Rather than being adopted solely on its own merits, DA and its variants became the primary mechanisms in England largely due to the significant failures of their main competitor—the first preference first system—based on criteria important enough to warrant changes to the School Admissions Code in England.
@@
A similar reform to those in BPS and England also took place in Chicago in 2009, albeit in a more abrupt manner. The city abandoned a capped version of the Boston mechanism used for the allocation of seats at selective high schools in the middle of the allocation process in 2009. A few months after the preferences were submitted under the capped Boston mechanism but before the placements were announced, officials at Chicago Public Schools (CPS) asked the students to resubmit their preferences under a capped version of DA. This incident was reported in a Chicago Sun-Times story as follows (Rossi, 2009):
@@
In addition to the manipulability of the capped version of the Boston mechanism, as highlighted by Huberman’s outrage, its failure to satisfy no justified envy was also a major concern in Chicago. Thus, in a redesign following a reform path similar to those in Boston and England, the Boston mechanism’s failure to meet two critical desiderata led to its replacement in yet another major school district, along with the adoption of a variant of DA. Remarkably, these failures were deemed so unacceptable that officials could not even wait for the next academic year to correct the mechanism, instead changing student assignments mid-process.
@@ 6.13 Broader Impact in Education Market Design
Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003), later reinforced by school choice reforms in New York City and Boston in the mid-2000s, marked the beginning of one of the most active and successful subfields of design economics: education market design. [^66] See Pathak (2017) and Abdulkadiroğlu and Andersson (2023) for excellent overviews of this literature. Since then, numerous school districts in the U.S. have adopted the DA mechanism, including Chicago Public Schools, Denver Public Schools, Indianapolis Public Schools, Newark Public Schools, the Recovery School District in New Orleans, Tulsa Public Schools, and Washington D.C. Public Schools (Pathak, 2017; Abdulkadiroğlu and Andersson, 2023). All these reforms have been shaped by design economists. To my knowledge, no jurisdiction reported using the DA mechanism in the U.S. or elsewhere prior to the publication of this literature. [^67] See Section 13.2 for Seattle’s unusual experience with the DA mechanism. Although the city replaced the Boston mechanism with the DA for its public schools in 1999, after Balinski and Sönmez (1999) but before Abdulkadiroğlu and Sönmez (2003) was published, it did not adequately communicate the reform to the public. It was not until 2008 that it became widely known that Seattle had been using the DA mechanism for nearly a decade. Once this information surfaced, the city reverted to the Boston mechanism within a few years.
@@
One possible explanation for why the DA mechanism and its variants have become the clear choice in higher education but not in primary or secondary education lies in differing attitudes toward the importance of the no justified envy axiom in these settings. As noted earlier in this section, Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003) argue that no justified envy is more likely to be viewed as indispensable in settings where priorities are determined through centralized exams. If this hypothesis holds, it could explain why versions of the DA are more prevalent in higher education, where reliance on performance metrics such as centralized exams is more common globally than in primary or secondary education.
As we have emphasized throughout this section, the vulnerability of the Boston mechanism to preference manipulation has been a key factor in its replacement by variants of the DA mechanism in many jurisdictions. Over the last two decades, numerous jurisdictions have also adopted hybrid mechanisms that combine features of the DA and Boston mechanisms to reduce vulnerability to preference manipulation. Among these, the most prominent is the parallel mechanism, first introduced in China in 2001 in Hunan province for military academies, later adopted by Jiangsu in 2005 and Zhejiang in 2007, and eventually implemented by 28 out of 31 provinces by 2012 (Chen and Kesten, 2017). [^68] See Section 13.3 for a closely related Taiwan Deduction mechanism.
@@
When there is a single band, the parallel mechanism becomes equivalent to the DA mechanism (or a capped version of DA), whereas when each band contains a single school, it becomes equivalent to the Boston mechanism. In many ways, the Chinese resolution to the Boston mechanism’s vulnerability to preference manipulation is not too different from that of jurisdictions such as New York City, which adopted capped versions of the DA mechanism rather than its strategy-proof unrestricted version.
@@ 7 Kidney Exchange
In the mid-2000s, through a series of papers, a team of design economists consisting of Alvin Roth, Utku Ünver, and me laid the theoretical foundation for organized kidney exchange (Roth et al., 2004, 2005b, 2007). Leveraging formal techniques from market design and optimization, kidney exchange rapidly transformed living donor kidney donation in numerous countries, saving thousands of lives annually within just a few years (Purtill, 2018; Rose, 2019). As an unexpected yet transformative application of market design, these developments played a pivotal role in elevating the field’s practical relevance and visibility during this period.
@@
But how did a team of economists manage to develop the tools and establish the infrastructure that now consistently touches so many lives? This section explores the pivotal role of key ideas in minimalist market design, illustrating their profound influence on the success and widespread adoption of organized kidney exchange clearinghouses worldwide. The policy and practical impacts of kidney exchange and school choice, in turn, later inspired my systematic application of these ideas in subsequent market design efforts and ultimately led to the conceptualization of minimalist market design, as presented in this monograph.
@@ 7.1 Paired Kidney Exchange and List Exchange
In the early 2000s, Ünver and I were both faculty members at Koç University in İstanbul. As discussed earlier in Section 5.6.2, our first joint project Sönmez and Ünver (2005) focused on the house allocation with existing tenants model (see Sections 5.4-5.6). Ünver spent the academic year 2002-2003 at Harvard University, visiting Roth. During this time, Roth mentioned to Ünver that this model has a direct healthcare application known as “kidney exchange.”
Naturally, discovering the connection between our work and a healthcare application naturally filled us with excitement and curiosity. We quickly learned that kidney exchange involves two key inventions by transplant surgeons aimed at improving the efficacy of living-donor kidney transplantation: paired kidney exchange and list exchange. Both inventions involve kidney patients and their willing living donors. However, due to immunological incompatibilities—either blood-type or tissue-type mismatch—a transplant from a donor to their paired patient may sometimes be infeasible or suboptimal. To address this barrier to living donation, transplantation experts developed these donor exchange arrangements, which were prominently discussed in a Consensus Statement on the Live Organ Donor, issued in 2000 by representatives of several professional transplantation societies (Abecassis et al., 2000). [^70] The Consensus Statement aimed to provide practice guidelines for transplant physicians, primary care providers, and healthcare planners.
@@
Paired Kidney Exchange: This invention involves an exchange of donors between two kidney patients, each of whom is typically incompatible with their own donor but compatible with the other patient’s donor. It is essentially an in-kind gift exchange between two patient-donor pairs to overcome biological incompatibility.
@@
This invention was originally proposed by a transplant surgeon in Rapaport (1986) and has been practiced in South Korea since the early 1990s (Park et al., 1999). It gained traction in the U.S. in the early 2000s, though only a handful of paired kidney exchanges had been conducted before our involvement in 2004. The biggest challenge was the absence of an organized system or even a database to facilitate these exchanges. At the time, methods for discovering paired kidney exchanges were limited to rudimentary practices, such as discussions between patients and their donors in settings like dialysis rooms.
@@
List Exchange: This invention, also known as an indirect exchange, was introduced in Ross and Woodle (2000) and involves exchanging a kidney from an incompatible living donor of a patient for a kidney from the deceased-donor (DD) list. The patient gains priority on the DD list if their incompatible donor instead donates a kidney to the DD list.
@@
Involving only a single patient-donor pair, list exchange is much easier to organize than paired kidney exchange. However, a major ethical concern, as highlighted by Ross and Woodle (2000), is its potential detrimental impact on blood-type O patients without living donors, whose options are limited to kidneys from the deceased donor (DD) list. Patient-donor pairs with a blood-type O donor are much less likely to participate in a list exchange than those with a blood-type O patient. [^71] There are four blood types: A, B, O, and AB. In the absence of other complications, most notably tissue-type incompatibility, a blood-type O kidney can be transplanted to any patient. A blood-type A kidney can be transplanted to patients with blood type A or AB. A blood-type B kidney can be transplanted to patients with blood type B or AB. A blood-type AB kidney can only be transplanted to patients with blood type AB.
@@
This disparity is particularly concerning in the U.S., where blood type O is disproportionately common among certain minority groups. Therefore, while appealing from a utilitarian perspective, list exchange raised significant concerns in terms of equity.
@@
Among these two inventions, paired kidney exchange garnered widespread approval in the Consensus Statement of the transplantation community, being deemed “ethically acceptable.” In contrast, in line with the equity concerns highlighted in Ross and Woodle (2000), the Consensus Statement underscored the ethical challenges associated with list exchange (Abecassis et al., 2000).

### v4

@@ 6.12 External Validity for Minimalist Market Design: School Choice Reforms in England and Chicago
In Section 1.2, we defined external validity as policy relevance that arises when field implementation aligns with earlier research, but importantly, when that alignment occurs independently as policymakers and other stakeholders converge on similar approaches and conclusions.
To illustrate, take the BPS school choice reform discussed in Section 6.10. A central trigger was vulnerability to preference manipulation: Boston officials abandoned the Boston mechanism after recognizing its manipulability and adopted a strategy-proof alternative, as recommended in Abdulkadiroğlu and Sönmez (2003). In his memo to the School Committee on May 25, 2005, Superintendent Thomas Payzant wrote (Abdulkadiroğlu et al., 2006):
To my knowledge, this was the first time strategy-proofness—a technical axiom from mechanism design—had been invoked by a policymaker in public deliberations to advocate a position. Moreover, Payzant framed strategy-proofness not in its usual mechanism-design role as an incentive-compatibility constraint, but as an equity principle. [^73] In economic theory, strategy-proofness is considered a special case of dominant-strategy incentive compatibility (DSIC), where private information typically takes the form of preferences over outcomes. As far as formal representation through mathematical relations is concerned, the two notions are indeed isomorphic. However, they have a subtle but very important distinction. DSIC is a concept in positive economics. In an effort to overcome challenges due to private information, it was formulated by Hurwicz (1960, 1972, 1973) as a constraint in his mechanism design framework. Thus, it mattered only to the extent that incentive considerations affected the outcome of an interaction between self-interested and rational economic agents. In particular, if the same optimum could be reached through dishonest revelation of private information, this framework would see no use for the constraint. It is a means to an end, not an end in itself. In contrast, strategy-proofness is both a positive and a normative concept. It is usually a key design objective for policymakers, serving as an end in itself. This represents direct policy impact, as Payzant’s framing drew explicitly on Abdulkadiroğlu and Sönmez (2003).
By contrast, when nearly identical reforms unfolded a few years later in England and Chicago, this time without guidance from economic designers, they reflected a different form of policy relevance. As we wrote in Pathak and Sönmez (2013):
@@
This notable parallel lends external validity to both the policy guidance of Abdulkadiroğlu and Sönmez (2003) and its interpretation of strategy-proofness as a key policy objective.
More importantly—and central to this monograph—it made clear that something larger was at play.
Beyond correcting flaws in original institutional designs, a central aim of minimalist market design is to recover their “intended” form—often by carving away incidental yet detrimental features that have accumulated over time. This makes the notion of external validity especially salient for research that follows this framework. Consequently, if decision-makers later (1) identify the same types of failures within their own institutions, (2) recognize these failures as stemming from the same underlying causes, and (3) independently devise resolutions that parallel those found in earlier research—effectively following the minimalist playbook on their own—this constitutes a particularly clear instance of external validity.
It was in observing precisely such instances—most strikingly in policymaker-driven reforms in England and Chicago—that the conceptual scope of this approach came into sharper focus. These reforms served as an intellectual awakening, revealing that what I had pursued largely by instinct was in fact part of a broader, coherent research and policy framework. With this heightened awareness, I began applying the minimalist framework more deliberately and systematically from the early 2010s—a shift catalyzed by the two cases in Pathak and Sönmez (2013), discussed next.
@@
In England, the nationwide 2003 School Admissions Code mandated that local authorities, an operating body similar to a U.S. school district, coordinate student admissions through a centralized mechanism. With the 2003 code, two classes of mechanisms were recommended.
The first of these two classes was DA and its “capped” versions, where the maximum number of schools that can be ranked in student preferences is limited to a fixed number. The second was a generalization of the Boston mechanism—called a first preference first system—and its capped versions. However, four years later, all versions of the first preference first system, including the Boston mechanism, were banned nationwide under the 2007 School Admissions Code. As a result, all local authorities in England were using variants of DA by 2007.
@@
Thus, much like in the 2005 reform at BPS, the Boston mechanism and its first preference first variant were abandoned in England due to their vulnerability to preference manipulation. Rather than being adopted solely on its own merits, DA emerged as the primary mechanism largely because of the critical weaknesses of its main competitor—failures serious enough to warrant changes to the School Admissions Code. The parallel is striking: the same flawed mechanism, the same root cause, and the same minimalist resolution.
This policymaker-driven reform provides a clear case of external validity for the research in Abdulkadiroğlu and Sönmez (2003) and demonstrates the power of minimalist market design to anticipate institutional evolution and shape enduring, policy-relevant reforms.
@@
A similar reform to those in BPS and England also took place in Chicago in 2009, albeit in a more abrupt manner. Midway through the allocation process, the city abandoned a capped version of the Boston mechanism that had been used for allocating seats at selective high schools. A few months after preferences were submitted under the capped Boston mechanism—but before placements were announced—officials at Chicago Public Schools (CPS) asked students to resubmit their preferences under a capped version of DA. This incident was reported in a Chicago Sun-Times story (Rossi, 2009):
@@
In addition to vulnerability to preference manipulation, the mechanism’s failure to satisfy no justified envy—evident from Huberman’s outrage—was a major concern in Chicago. As in Boston and England, the same failure and the same root cause again led to the same solution: replacement of a capped version of the Boston mechanism with a variant of DA. These failures were deemed so unacceptable that officials did not wait for the next academic year, instead altering student assignments mid-process.
This episode provides yet another compelling example of external validity for the research in Abdulkadiroğlu and Sönmez (2003) and underscores the broader potential of minimalist market design to generate policy-relevant insights.
@@ 6.13 Broader Impact in Education Market Design
Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003), later reinforced by school choice reforms in New York City and Boston in the mid-2000s, marked the beginning of one of the most active and successful subfields of economic design: education market design. [^74] See Pathak (2017) and Abdulkadiroğlu and Andersson (2023) for excellent overviews of this literature. Since then, numerous school districts in the U.S. have adopted the DA mechanism, including Chicago Public Schools, Denver Public Schools, Indianapolis Public Schools, Newark Public Schools, the Recovery School District in New Orleans, Tulsa Public Schools, and Washington D.C. Public Schools (Pathak, 2017; Abdulkadiroğlu and Andersson, 2023). All these reforms have been shaped by economic designers. To my knowledge, no jurisdiction reported using the DA mechanism in the U.S. or elsewhere prior to the publication of this literature. [^75] See Section 13.2 for Seattle’s unusual experience with the DA mechanism. Although the city replaced the Boston mechanism with the DA for its public schools in 1999, after Balinski and Sönmez (1999) but before Abdulkadiroğlu and Sönmez (2003) was published, it did not adequately communicate the reform to the public. It was not until 2008 that it became widely known that Seattle had been using the DA mechanism for nearly a decade. Once this information surfaced, the city reverted to the Boston mechanism within a few years.
@@
One possible explanation for why the DA mechanism and its variants have become the clear choice in higher education but not in primary or secondary education lies in differing attitudes toward the importance of the no justified envy axiom in these settings. As noted earlier in this section, Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003) argue that no justified envy is more likely to be regarded as indispensable in settings where priorities are determined through centralized exams. If this hypothesis holds, it could explain why variants of DA are more prevalent in higher education, where reliance on performance metrics such as centralized exams is more common globally than in primary or secondary education. [^76] See the Rossi (2009) quote in Section 6.12.2 regarding the school choice crisis in Chicago Public Schools for evidence that lends support to this hypothesis.
As we have emphasized throughout this section, the vulnerability of the Boston mechanism to preference manipulation has been a key factor in its replacement by variants of the DA mechanism in many jurisdictions. Along these lines, but as a partial compromise, numerous jurisdictions have also adopted hybrid mechanisms over the last two decades that combine features of the DA and Boston mechanisms to reduce vulnerability to preference manipulation. Among these, perhaps the most prominent is the parallel mechanism, first introduced in China in 2001 in Hunan province for military academies, later adopted by Jiangsu in 2005 and Zhejiang in 2007, and eventually implemented by 28 out of 31 provinces by 2012 (Chen and Kesten, 2017). [^77] See Section 13.3 for a closely related Taiwan Deduction mechanism.
@@
When there is a single band, the parallel mechanism reduces to the DA mechanism (or a capped version), and when each band contains only one school, it reduces to the Boston mechanism (or a capped version), illustrating how the parallel mechanism functions as a hybrid of these two designs.
@@ 7 Kidney Exchange
In the mid-2000s, through a series of papers, a team of economic designers consisting of Alvin Roth, Utku Ünver, and me laid the theoretical foundation for organized kidney exchange (Roth et al., 2004, 2005b, 2007). Leveraging formal techniques from market design and optimization, kidney exchange rapidly transformed living donor kidney donation in numerous countries, saving thousands of lives annually within just a few years (Purtill, 2018; Rose, 2019). As an unexpected yet transformative application of market design, these developments played a pivotal role in elevating the field’s practical relevance and visibility during this period.
@@
But how did a team of economists manage to develop the tools and establish the infrastructure that now consistently touches so many lives? This section explores the pivotal role of key ideas in minimalist market design, illustrating their profound influence on the success and widespread adoption of organized kidney exchange clearinghouses worldwide. The policy and practical impacts of kidney exchange and school choice in the late 2000s—amplified by the external validity demonstrated by school choice reforms in England and Chicago in the early 2010s—in turn inspired my more systematic application of these ideas in subsequent market design efforts, ultimately leading to the conceptualization of minimalist market design as presented in this monograph.
@@ 7.1 Paired Kidney Exchange and List Exchange
End-Stage Renal Disease (ESRD) is a major global health problem. As of 2025, more than 800,000 people in the U.S. alone—about 0.2% of the population—are living with ESRD. Kidney transplantation is the preferred treatment, but kidneys from deceased donors fall well short of the need, making living donors a crucial source of transplant kidneys.
Unfortunately, immunological barriers can make transplantation between a living donor and their intended recipient infeasible or medically suboptimal. One of the most consequential barriers is ABO blood group incompatibility.
There are four main human blood types—A blood type, B blood type, AB blood type, and O blood type—which are defined by the presence or absence of two antigens on red blood cells, A and B. Individuals with type O have neither antigen; [^79] The term “O” comes from the German word Ohne, meaning “without.” In many German-speaking regions, blood type O is still referred to as Blutgruppe 0 (“blood group zero”). those with type A have only the A antigen, those with type B have only the B antigen, and those with type AB have both.
Figure 15: Directed graph representation of blood-type compatibility under the ABO grouping.
People naturally produce antibodies against the antigens they lack. Those without the A antigen produce anti-A antibodies, and those without the B antigen produce anti-B antibodies. As a result, someone lacking the A antigen cannot safely receive blood or a solid organ from someone who has it, and likewise for the B antigen.
This creates a major biological asymmetry that puts type O patients at a clear disadvantage: because they lack both antigens, they can receive kidneys only from type O donors. More broadly, in the absence of other complications—most notably tissue-type incompatibility—a kidney from a type O donor can be transplanted into any patient; a kidney from a type A donor into patients with type A or AB; a kidney from a type B donor into patients with type B or AB; and a kidney from a type AB donor only into patients with type AB. [^80] One exception, discouraged in much of the world, is ABO-incompatible kidney transplantation, which uses intensive immunosuppression to overcome blood-type incompatibility. It has been adopted in a few countries such as Japan and South Korea, where cultural or historical factors have severely limited the availability of deceased-donor kidneys, but is generally not used in most other countries facing similar or even greater deceased-donor shortages, such as Turkey, where it is not covered by national health insurance. Because it often yields inferior outcomes compared to compatible transplants, it is typically avoided due to the extensive preparation it requires and its higher medical risks and costs. Figure 15 illustrates these compatibility relations as a directed graph under the ABO grouping.
This structural disadvantage for patients with blood type O has long raised ethical and policy concerns within the transplant community, prompting efforts to develop allocation practices that could mitigate such inequities.
In 2000, representatives from several professional transplantation societies issued the Consensus Statement on the Live Organ Donor (Abecassis et al., 2000), aiming to provide standardized ethical, medical, and procedural guidelines for transplant physicians, primary care providers, and healthcare planners. Central to this discussion, the statement discussed two types of donor exchange arrangements to overcome barriers to living donation—paired kidney exchange and list exchange—with strong enthusiasm for the first and caution advised on the second.
Paired Kidney Exchange.
@@
This innovation involves an exchange of donors between two kidney patients, each of whom is (typically) incompatible with their own donor but compatible with the other patient’s donor (see Figure 16 for an illustration). It is essentially an in-kind gift exchange between two patient–donor pairs to overcome biological incompatibility.
Figure 16: Paired kidney exchange. Patient P1 is incompatible with the kidney K1 from their own living donor but is compatible with the kidney K2 from Patient P2’s living donor. Similarly, Patient P2 is incompatible with the kidney K2 from their own living donor but is compatible with the kidney K1 from Patient P1’s living donor. By exchanging donors, both patients can receive compatible transplants.
@@
Originally proposed by Rapaport (1986)—a transplant surgeon—this approach has been practiced in South Korea since the early 1990s (Park et al., 1999). [^81] South Korea was the first country in the world to implement kidney exchange, launching a program at Hanyang University Hospital in 1991 (Park et al., 1999). Between 1991–2010, 152 kidney exchange transplants were performed at this center (Kim et al., 2012). Another leading center at Yonsei University reported 129 kidney exchange transplants between 1995–2006 (Huh et al., 2008). However, with the introduction of desensitization protocols and ABO-incompatible living-donor transplantation, kidney exchange declined, and by the 2010s only a handful of operations were performed annually in the country. It gained traction in the U.S. in the early 2000s, though only a handful of paired kidney exchanges had been conducted before our involvement in 2004. The biggest challenge was the absence of an organized system or even a database to facilitate these exchanges. At the time, methods for discovering paired kidney exchanges were limited to rudimentary practices, such as discussions between patients and their donors in settings like dialysis rooms.
List Exchange.
@@
This innovation, also known as an indirect exchange, involves a donor giving a kidney to the deceased-donor (DD) list in exchange for their incompatible patient receiving priority on the DD list (see Figure 17 for an illustration).
@@
In their paper introducing this approach, Ross and Woodle (2000) also highlighted a major ethical concern: its potential detrimental impact on blood-type O patients without living donors, whose options are limited to kidneys from the DD list. Involving only a single patient–donor pair, list exchange is much easier to organize than paired kidney exchange, but pairs with a blood-type O donor are far less likely to participate than those with a blood-type O patient.
@@
This disparity is particularly concerning in the U.S., where blood type O is disproportionately common among certain minority groups. Therefore, while appealing from a utilitarian perspective, list exchange raises significant concerns in terms of equity.
Figure 17: List exchange. Patient P1 is incompatible with kidney K1 from their living donor. The donor instead donates their kidney K1 to the deceased-donor list, where it is allocated to the highest-priority patient L. In return, Patient P1 receives priority on the deceased-donor list and is allocated a deceased-donor kidney D.
@@
Among these two approaches, paired kidney exchange garnered widespread approval in the Consensus Statement on the Live Organ Donor (Abecassis et al., 2000), being deemed “ethically acceptable.” In contrast, in line with the equity concerns highlighted in Ross and Woodle (2000), the Consensus Statement underscored the ethical challenges associated with list exchange.